OpenStudy (anonymous):

A rectangular Persian carpet has a perimeter of 18 inches. The length of the carpet is 22 inches more than the width. What are dimensions of the carpet?

OpenStudy (kainui):

Try to write an equation based on this. Draw a picture and label stuff like L for length and W for width. Take a guess and I'll help.

OpenStudy (anonymous):

okay, would it be 22w+L= 18?

OpenStudy (ranga):

Length of the carpet is 22 inches more than the width means: L = W + 22 ------ equation (1) That is one equation. Another equation is Perimeter = 18 That means: 2(L + W) = 18 ------ equation (2) You have 2 equations and 2 unknowns. Solve for L & W.

OpenStudy (kainui):

Almost micheller, so if the length is 22 more than the width, you can say: L is length "is" means equals 22 is of course just the number more means add W is width so L=22+W. Now you should be able to see what the perimeter is based on length and width by looking at the picture and adding up all the sides.

OpenStudy (anonymous):

@Kainui is the equation 2(L+W)=18

OpenStudy (kainui):

Yes exactly! So now you have two equations and two unknowns. That means you can solve for both of your variables now. =D

OpenStudy (anonymous):

Thanks